OpenStudy (anonymous):

A man 5 feet tall walks at the rate 4 feet per second directly away from a street light wich is 20 feet above the street. B) At what rate is the length of his shadow changing? A) Is the length increasing or decreasing?

OpenStudy (amistre64):



OpenStudy (amistre64):

hmm

OpenStudy (amistre64):

we might be able to use similar triangles to compare with

OpenStudy (amistre64):



OpenStudy (amistre64):

\[\frac AB=\frac ab\] B = walking + b, and lets change b to s for shadow \[\frac A{w+s}=\frac as\] \[As=a({w+s})\] \[A's+As'=a'({w+s})+a({w+s})'\] since A and a are constants, their derivatives go to zero \[As'=a(w+s)'\] \[As'=aw'+as'\] \[As'-as'=aw'\] \[s'(A-a)=aw'\] \[s'=\frac{aw'}{(A-a)}\] so with any luck \[s'=\frac{5(4)}{(20-5)}\]

OpenStudy (anonymous):

Thank you so much! That helped a lot.

OpenStudy (amistre64):

youre welcome, and with any luck it was right :) is there a way to check?

OpenStudy (anonymous):

Yes it was right. I have the answer but I didn't know how to solve it. Thanks :)